OpenStudy (anonymous):

One number is eight more than twice another. If their difference is 25, what is the larger number? a)17 b)42 c)35

OpenStudy (radar):

x= one number, y = the other number, x=8+2y x-y=25. You have two equations with two unknowns. You can solve. 8+2y-y=25 (substituting for x) solving: y=25-8=17, so x=42 from x-17=25 x is the larger number and is 42